OpenStudy (anonymous):

Stuck: trying to solve for x: cos(2x)=-sin(x) where 0<=x<=2pi My work below.

OpenStudy (anonymous):

\[\cos^{2}x-\sin^{2}x=-\sin(x)\]

OpenStudy (anonymous):

After that, a bunch of reshuffling that got me nowhere.

OpenStudy (anonymous):

If you use \[\cos 2x = 1 - \sin ^{2} x\] you will have a quadratic in sin x

OpenStudy (anonymous):

sorry, 1 - 2(sin x)^2

OpenStudy (anonymous):

So, you use \[\cos 2x = 1 - 2\sin ^{2}x\]

OpenStudy (anonymous):

cos(2x) is equal to 1 - 2sin^2(x) This makes it all 1-2sin^2(x) = -sin(x) re-arrange to make 2sin^2(x) - sin(x) - 1 =0

OpenStudy (anonymous):

(2sinx-1)(sinx+1) sinx = -1 or sinx= 0.5

OpenStudy (anonymous):

Would it have been easier if I had given the original problem? \[\cos(2x)=-\sin(x)\]

OpenStudy (anonymous):

that's the same problem as your initial question, i was able to see what you mean

OpenStudy (anonymous):

Oh geez, ignore. SO tired. Four hours sleep. LOL!

OpenStudy (anonymous):

:D i know the feeling